# Prediction of the Most Fire-Sensitive Point in Building Structures with Differentiable Agents for Thermal Simulators

### v1

@@ VII-A2 Evaluation metrics
We evaluate the performance of the MIDR Predictor using three metrics: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Spearman’s rank correlation coefficient. MSE, defined in Equation (17), quantifies the squared difference between predicted and ground truth MIDRs. MAE measures the absolute difference and is defined as:
$L_{\text{MAE}}=\frac{1}{\sum_{m=1}^{M}N_{m}}\sum_{m=1}^{M}\sum_{n=1}^{N_{m}}\left|\widehat{d}_{m,n,\max}-d_{m,n,\max}^{\,\text{gt}}\right|,$ (20)
where $d_{m,n,\max}^{\,\text{gt}}$ is the ground truth MIDR of $m$-th structure under the $n$-th fire scenario (calculated by OpenSeesRT), and $\widehat{d}_{m,n,\max}$ is the corresponding prediction from the MIDR Predictor.
Since the ultimate goal is to identify the MFSP rather than exact MIDR values, the ranking of MIDRs across different fire scenarios for each structure is critical. For this, we use Spearman’s rank correlation coefficient ($\rho_{s}$), which measures the monotonic relationship between two sequences $X$ and $Y$ and is defined as follows:
$\rho_{s}\left(X,Y\right)\triangleq\rho\left(R_{X},R_{Y}\right)=\frac{\text{cov}\left(R_{X},R_{Y}\right)}{\sigma_{R_{X}}\sigma_{R_{Y}}}.$ (21)
In this formulation, $R_{X}$ and $R_{Y}$ represent the rank-ordered sequences of $X$ and $Y$, respectively, where each element is replaced by its rank in the sorted sequence. Moreover, $\rho(\cdot,\cdot)$ denotes the Pearson correlation coefficient, $\text{cov}\left(R_{X},R_{Y}\right)$ represents the covariance of $R_{X}$ and $R_{Y}$, which quantifies how changes in rank of $X$ correspond to changes in rank of $Y$, and $\sigma_{R_{X}}$ and $\sigma_{R_{Y}}$ denote the standard deviations of the rank distributions $R_{X}$ and $R_{Y}$, respectively. Unlike Pearson’s correlation, which measures linear relationships, Spearman’s correlation focuses on how well the rank-ordering of the variables aligns. This makes it particularly suitable for ordinal data or nonlinear relationships. The resulting values of $\rho_{s}$ range from $-1$ to $+1$, with values closer to $+1$ indicating strong agreement in rank-ordering, values near 00 indicating weak or no correlation, and values closer to $-1$ indicating a strong inverse correlation.
In this paper, for the $m$-th structure with $N_{m}$ fire scenarios, we compute $\rho_{s}\left(X,Y\right)$ between the sequence of ground truth results, $X\equiv\left\{d_{m,1,\max}^{\,\text{gt}},d_{m,2,\max}^{\,\text{gt}},\ldots,d_{m,N_{m},\max}^{\,\text{gt}}\right\}$, obtained from OpenSeesRT simulations and the sequence of predictions, $Y\equiv\left\{\widehat{d}_{m,1,\max},\widehat{d}_{m,2,\max},\ldots,\widehat{d}_{m,N_{m},\max}\right\}$, produced by the MIDR Predictor. The final evaluation metric is the average $\rho_{s}$ across all structures, providing a robust measure of prediction quality for MFSP identification.
@@ VII-A3 Evaluation results
For training and evaluating the MIDR Predictor, we exclusively use the labeled dataset, splitting it into $80\%$ for training and $20\%$ for testing. Various neural network configurations and training methods were compared, with the results summarized in Table V.
@@
In Table V, the methods Strawman 1, Strawman 2, and Proposed are described in Section V-B and illustrated in Figure 7. For Strawman 2 and Proposed methods, we further assess the impact of enabling the EU functionality. Two key observations are summarized below:
  • EU functionality: Comparisons between cases with and without EU reveal that enabling EU consistently enhances model performance. Specifically, the inclusion of EU increases $\rho_{s}$ in all cases, and reduces both MAE and MSE for not only the proposed predictors but also the Strawman 2 predictor based on small network.
@@
Figure 9(a) presents the Complementary Cumulative Distribution Function (CCDF) of $\rho_{s}$, i.e., ¯${\overline{{\mathcal{F}}}}_{\rho_{s}}(a)={\mathcal{P}}(\rho_{s}\geq a)$, where $\mathcal{P}$ indicates the probability, for the proposed method with EU. Key observations include:
  • For predictors using small and large networks (solid lines), namely small and large predictors for simplicity, $83.2\%$ of structures achieve a $\rho_{s}$ value greater than 0.50, and $64.4\%$ (small predictor) and $66.7\%$ (large predictor) of structures achieve a $\rho_{s}$ value exceeding $0.75$. Additionally, large predictor improves on the worst case, where the smallest $\rho_{s}$ value increases from $-0.86$ for small predictor to $-0.58$ for large predictor.
  • MIDR predictors perform much better on severe cases (dashed lines), defined for a structure as a condition where the ground truth MIDR exceeds $2\%$ under at least one fire case. This is in comparison with all cases (solid lines). In severe cases, for the small predictor, $98.0\%$ of structures achieve a $\rho_{s}$ value greater than $0.5$, and $89.3\%$ exceed $0.75$ with an average of $0.90$. For the large predictor, $98.7\%$ of structures achieve a $\rho_{s}$ value greater than $0.5$, while $90.0\%$ exceed $0.75$ with an average of $0.91$.
These results indicate that the proposed method performs exceptionally well under severe cases. This is consistent with our expectations: when building structures are less sensitive to fire, errors in MIDR prediction have a relatively small impact. However, the risk of MIDR prediction errors becomes critical when structures are sensitive to fire.
To provide a more intuitive understanding of the proposed method’s performance, we present the ground truth MIDR and predicted MIDR for three example structures under the small predictor (Figure 9(b)) and the large predictor (Figure 9(c)). In these examples, ex1 and ex2 represent severe cases, while ex3 remains relatively stable with low MIDR across all fire scenarios. From these figures, for both predictors, the predicted results for ex1 and ex2 closely match the ground truth, achieving $\rho_{s}$ values above $0.92$. For ex3, the performance of both predictors is lower but the large predictor outperforms the small predictor, achieving a $\rho_{s}$ value above $0.7$.
@@
Figure 9: Representative MIDR Predictor results: (a) CDF of $\rho_{s}$ of proposed MIDR Predictors for two different sizes (severe cases for structures whose ground truth MIDR is larger than $2\%$ in at least one fire case); (b) and (c) show the ground truth and predicted MIDR of 3 example structures (ex1, ex2 & ex3) for the small and large predictors, respectively.
@@
In conclusion, the proposed method with EU functionality demonstrates exceptional performance across all metrics, particularly for severe cases, where accurate MIDR prediction is most critical. These findings highlight the robustness and effectiveness of the proposed method in supporting fire safety assessments for building structures.
@@ VII-B MFSP Predictor Evaluation
To evaluate the MFSP Predictor, we adopt the proposed method with EU as the final model for the MIDR Predictor and utilize it to train the MFSP Predictor. As described in Section VI-B, we consider two training strategies for the MFSP Predictor:
  1. 1.
    De novo training: Training a completely new network from scratch.
  2. 2.
    TL with GNN: Reusing the GNN module, including the initial encoders, from the MIDR Predictor.
@@
  1. 1.
    MSE loss: Mean squared error focusing on accuracy of predicted coordinates.
  2. 2.
    Hybrid loss: A combination of MSE loss and the MIDR-based loss for better ranking consistency.
This section compares the performance of these four combinations. The network architecture of the MFSP Predictor is almost identical to that of the MIDR Predictor, consisting of an initial encoder, a GNN, pooling layers, and an MLP task head. There are only a few differences:
  • Final MLP Head: The task head in MFSP Predictor replaces the linear layer in the MIDR Predictor with a single-hidden-layer MLP. The hidden layer’s dimension matches the graph embedding (twice the node embedding dimension), and the output layer has a dimension of 3, representing the predicted coordinates $(x,y,z)$. A “sigmoid” activation function is applied to normalize the predicted coordinates to the range $(0,1)$, aligning with the building normalized spatial dimensions.
  • De Novo training: The initial node input features include only the spatial coordinates of the nodes. Fire point information is excluded, as no prior knowledge from the MIDR Predictor is utilized.
  • TL with GNN training: The GNN module and initial encoders from the MIDR Predictor are directly reused. Moreover, the fire point information is replaced with a VFP to ensure consistency in input structure. This enables the network to generalize its understanding of structural information without relying on specific fire locations.
By evaluating these configurations and loss functions, the MFSP Predictor aims to efficiently and accurately predict the MFSP within the building.
@@ VII-B1 Evaluation metrics
In Section VI-A, we introduce the “Pseudo-Labeling” method, which generates pseudo ground truth coordinates for the MFSP. These pseudo-labels are used to evaluate the MFSP Predictor through various metrics, capturing both point-level and room-level performances.
For a given $m$-th structure, the distance error between predicted MFSP and pseudo ground truth MFSP is defined as:
@@
where $\widehat{x}_{m}$, $\widehat{y}_{m}$, $\widehat{z}_{m}$ represent the $x$, $y$, $z$ coordinates predicted by the MFSP Predictor, while $x_{m}^{\,\text{pgt}}$, $y_{m}^{\,\text{pgt}}$, $z_{m}^{\,\text{pgt}}$ denote the pseudo ground truth MFSP coordinates obtained using the MIDR Predictor. This metric provides a direct measure of the prediction accuracy at the point level. However, as the pseudo ground truth is defined at a room-level granularity, its reliability may vary depending on the building’s size and complexity.
To align the evaluation with practicality, two room-level metrics are introduced: room distance error and room rank. These metrics treat rooms as the basic unit of measurement, reflecting how well the MFSP Predictor identifies the MFSR, defined as the room where MFSP is located.
Room distance error measures the distance between the predicted MFSR and the pseudo ground truth MFSR, assuming all edges (beams and columns) of the building structure have unit length. For the $m$-th structure, the room dimensions are defined as $\tilde{x}_{m}$, $\tilde{y}_{m}$, $\tilde{z}_{m}$ for the respective length, width, and height of each room. The coordinates of the center point of the predicted MFSR is denoted as $\left(\widehat{x}_{m}^{\prime},\widehat{y}_{m}^{\prime},\widehat{z}_{m}^{\prime}\right)$. The room distance is then computed as:
@@
This metric provides a scaled distance measure, accounting for the room size and structure dimensions.
Room rank evaluates the ranking of the room containing the predicted MFSP based on pseudo-labeled MIDR values. For the $m$-th structure with $N_{m}$ rooms (i.e., $N_{m}$ pseudo fire cases), the pseudo-labeled MIDRs are sorted in descending order, indexed by the fire case occurring at each room’s center point, i.e., $d_{m,1,\max}^{\,\text{pgt}}>d_{m,2,\max}^{\,\text{pgt}}>\cdots>d_{m,N_{m},\max}^{\,\text{pgt}}$. If the predicted MFSP falls into the $l$-th ranked room, the room rank is defined as $r_{m}=l$. Lower ranks indicate better prediction performance, as the predicted MFSR is closer to the MFSP.
In addition to the distance error, room distance error, and room rank, the predicted MIDR at the identified MFSP can also serve as an evaluation metric. Higher MIDR values reflect the Predictor’s ability to identify the MFSPs, directly correlating with the severity of the structural response.
These evaluation metrics collectively assess the MFSP Predictor’s performance, balancing point-level precision, room-level granularity, and the functional implications of predictions. The inclusion of room distance error and room rank ensures that the evaluation aligns with practical building-level assessments, where rooms are typically used reference units for structural analysis and fire safety planning.
VII-B2 Evaluation results
Using the metrics defined in the previous section, we evaluate the performance of the MFSP Predictor by calculating average quantities across all structures. Since MIDR Predictor provides a computationally efficient agent model to rapidly generate pseudo labels for unlabeled data, the large unlabeled dataset can be effectively utilized for training the MFSP Predictor. For this evaluation, the unlabeled dataset is divided into $80\%$ for training and $20\%$ for testing. The evaluation results for various training methods and loss functions are summarized in Table VI. For the “Hybrid loss”, the weight for the MSE term is fixed at $w_{2}=1\,\mathrm{m}^{-2}$, while the weight for the MIDR term is varied as $w_{1}=10$, $50$, or $100$, as described in Equation (18). In Table VI, the TL with GNN results for $w_{1}=50$, using both small and large networks, are taken as the reference cases for the listed multipliers in parentheses.
TABLE VI: Evaluation results of MFSP Predictor for different training methods and loss functions (numbers in boldface indicate best results).

### v5

@@ VII-A2 Evaluation metrics
We evaluate the performance of the MIDR predictor using three metrics: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Spearman’s rank correlation coefficient. MSE and MAE are respectively defined as follows:
$\text{MSE}=\frac{1}{\sum_{m=1}^{M}N_{m}}\sum_{m=1}^{M}\sum_{n=1}^{N_{m}}\left(\widehat{d}_{m,n,\max}-d_{m,n,\max}^{\,\text{gt}}\right)^{2},$ (20)
$\text{MAE}=\frac{1}{\sum_{m=1}^{M}N_{m}}\sum_{m=1}^{M}\sum_{n=1}^{N_{m}}\left|\widehat{d}_{m,n,\max}-d_{m,n,\max}^{\,\text{gt}}\right|,$ (21)
where $d_{m,n,\max}^{\,\text{gt}}$ is the ground truth MIDR of structure $m$ under fire scenario $n$, obtained from $\chi$ara, and $\widehat{d}_{m,n,\max}$ is the corresponding prediction from the MIDR predictor.
Since our ultimate goal is to train an argmaxer to identify the MFSP, rather than merely predicting the MIDR values, accurately assessing the relative ranking of MIDRs across different fire scenarios for each structure is critical. MSE and MAE can be good metrics for characterizing the closeness between the predicted MIDRs and the ground truth MIDRs. However, they do not adequately capture the ranking performance. Therefore, we introduce the Spearman’s rank correlation coefficient, $\rho_{s}$, which measures the monotonic relationship between two random variables. Formally, let $A$ and $B$ be two random variables, and let $R_{A}$ and $R_{B}$ denote their corresponding rank-transformed values based on the observed samples. The Spearman’s rank correlation coefficient is defined as:
$\rho_{s}\left(A,B\right)\triangleq\rho\left(R_{A},R_{B}\right)=\frac{\text{cov}\left(R_{A},R_{B}\right)}{\sigma_{R_{A}}\sigma_{R_{B}}},$ (22)
where $\rho(\cdot,\cdot)$ is the Pearson correlation coefficient, $\text{cov}\left(R_{A},R_{B}\right)$ is the covariance between the ranked values, and $\sigma_{R_{A}}$ $\sigma_{R_{B}}$ denote their respective standard deviations. Unlike Pearson’s correlation, which measures linear relationships, Spearman’s correlation assesses how well the orderings of the data match, making it especially suitable for ordinal data or for relationships that are monotonic but not necessarily linear. The range of $\rho_{s}$ is $[-1,+1]$, with values close to $+1$, 00, and $-1$ respectively indicating strong positive, weak or no, and strong negative monotonic associations.
In this paper, for structure $m$ with $N_{m}$ fire scenarios, we treat the ground truth and predicted MIDRs as realizations of underlying random variables. Specifically, we define the ground truth sample as $A\equiv\left\{d_{m,1,\max}^{\,\text{gt}},d_{m,2,\max}^{\,\text{gt}},\ldots,d_{m,N_{m},\max}^{\,\text{gt}}\right\}$, obtained from $\chi$ara simulations and the prediction samples as $B\equiv\left\{\widehat{d}_{m,1,\max},\widehat{d}_{m,2,\max},\ldots,\widehat{d}_{m,N_{m},\max}\right\}$, produced by the MIDR predictor. The final evaluation metric is the average $\rho_{s}$ across all structures, providing a robust measure of the prediction quality for the MFSP identification.
@@ VII-A3 Evaluation results
For training and evaluating the MIDR predictor, we exclusively use the labeled dataset, splitting it into $80\%$ for training and $20\%$ for testing. Various GNN configurations and training methods are compared, with the results summarized in Table V.
@@
Overall accuracy performance analysis
In Table V, the methods Strawman 1, Strawman 2, and Proposed are described in Section IV-B and illustrated in Figure 3. For Strawman 2 and Proposed methods, we further assess the impact of enabling the EU functionality. Two key observations are summarized below:
  • EU functionality: Comparisons between cases with and without EU reveal that enabling EU consistently enhances model performance. Specifically, the inclusion of EU increases $\rho_{s}$ in all cases, and reduces both MAE and MSE for not only the proposed predictor but also the Strawman 2 predictor based on small network.
@@
In structure design, for fire-sensitive buildings, it is expected that the location of MFSP should be predicted more accurately. On the contrary, if a building is not likely to collapse no matter where a fire occurs, then the tolerance for the MFSP prediction errors is higher. For this reason, we specifically consider those buildings whose ground truth MIDR exceeds $2\%$ in at least one fire case, and denote their data as severe cases.
We divide the data into severe cases and non-severe cases. In the former, the average variance of the MIDR of each building in the 30 fire scenarios reaches $0.40$, while this value is only $0.02$ in the latter. In other words, for severe cases, MIDRs vary greatly with the location of the fire points. In such cases, $\rho_{s}$ is an important indicator to measure the reliability of using the MIDR predictor to train the MFSP predictor, and results show that the average $\rho_{s}$ for severe cases reaches over $0.90$. Figure 9(a) presents the Complementary Cumulative Distribution Function (CCDF) of $\rho_{s}$, i.e., ¯${\overline{{\cal{F}}}}_{\rho_{s}}(a)={\cal{P}}(\rho_{s}\geq a)$, where $\cal{P}$ indicates the probability, for the proposed method with EU. The solid lines include both severe and non-severe cases, i.e., all the test data, where $83.2\%$ of structures achieve $\rho_{s}>0.50$, and using small and large predictors $64.4\%$ and $66.7\%$ of structures, respectively, achieve $\rho_{s}>0.75$. Additionally, the large predictor improves on the worst case, where the smallest $\rho_{s}$ increases from $-0.86$ for the small predictor to $-0.58$ for the large predictor. On the other hand, the dashed lines are for the severe cases where for the small predictor, $98.0\%$ of structures achieve $\rho_{s}>0.50$ and $89.3\%$ exceed $0.75$ with an average of $0.90$. For the large predictor, $98.7\%$ of structures achieve $\rho_{s}>0.50$, while $90.0\%$ exceed $0.75$ with an average of $0.91$.
For the non-severe cases, low MSE of predicted MIDRs guarantees the reliability of the MIDR predictors. Results show that the small predictor leads to an MSE as low as $0.06$ and the value for the large predictor is $0.05$. In the following, the differences between severe and non-severe cases are discussed in detail with three example structures.
(a) CCDF results ($\rho_{s}$-axis inverted).
(b) Small predictor.
(c) Large predictor.
@@
Figure 9: Representative MIDR predictor results: (a) CDF of $\rho_{s}$ for the proposed MIDR predictors for two different sizes (severe cases are for structures whose ground truth MIDR $>2\%$ in at least one fire case); (b) and (c) show the MIDR predictions vs. the ground truth for 3 example structures (ex1, ex2 & ex3) for the small and large predictors, respectively.
Exemplary illustration & time efficiency
Here, we present the predicted MIDR vs. the ground truth MIDR for three example structures (i.e., ex1, ex2 & ex3 in Figure 6) under the small (Figure 9(b)) and large (Figure 9(c)) predictors. In these examples, ex1 and ex2 represent severe cases, while ex3 remains relatively stable with low MIDR across all fire scenarios, i.e., a non-severe case. As shown in the figures, the predicted results for ex1 and ex2 closely match the ground truth for both predictors, achieving $\rho_{s}>0.92$. For ex3, the $\rho_{s}$ of both predictors is lower, but the large predictor outperforms the small one, achieving $\rho_{s}>0.70$. Although $\rho_{s}$ for ex3 is relatively low, the performance remains accurate, as the predicted and ground truth MIDR values are closely aligned along the ideal $45^{\circ}$ line with and the MSE $\leq 0.01$. Note that Spearman’s rank correlation coefficient does not evaluate the closeness between predicted and ground truth values, but rather evaluates the monotonic relationship. Therefore, in the consequent task of training an argmaxer, the high value of average $\rho_{s}\geq 0.90$ guarantees the reliability and validity. Moreover, for the non-severe cases, the low MSE indicates that the MIDR predictor is a good surrogate model of thermal simulators.
The differentiable agent shows great improvement on the time efficiency compared with traditional FEA. On our experimental machine, with 2 cores of Skylake CPU (2.1 GHz) and one GTX2080TI GPU, the average time consumptions for predicting the MIDR for the 3 example structures is listed in Table VI, where the number of nodes and edges are also listed for reference. Note that the values for $\chi$ara are the time consumptions for just one fire scenario, and the values for the agent are those for all 30 fire scenarios. We clearly see that the conventional FEA simulation’s efficiency is highly affected by the number of nodes and edges, while these numbers have less impact on the agent’s time consumptions. The most important finding, as expected, is that the GNN-based agent is much faster than the FEA simulations, with more than 3 orders of magnitude difference, i.e., $\chi$ara costs several to tens of seconds for only one fire scenario, while the agents cost only several milliseconds (ms) for a batch of 30 fire scenarios. Note that to obtain accurate time consumption of the agent, we run the agent for 1,000 times and report their average run time.
TABLE VI: Average time consumptions of different methods to obtain the MIDR for 3 examples in Figures 9(b) & (c).
ex Number of $\chi$ara Agent (1 batch of 30 fires)
Nodes Edges (1 fire) (ms) Small (ms) Large (ms)
1 331 728 70,950 8.63 12.64
2 286 618 48,340 8.59 11.00
3 59 101 6,590 4.12 4.29
@@
In conclusion, the proposed method with EU functionality demonstrates exceptional performance across all metrics for both accuracy and time efficiency, particularly for severe cases. These findings highlight the robustness and effectiveness of the method in supporting fire safety assessments for building structures. The high values of $\rho_{s}$, and the extremely low time costs also demonstrate the reliability and the efficiency of using the MIDR predictor as an agent for pseudo-labeling and as a direct indicator for training the MFSP predictor.
@@ VII-B MFSP Predictor Evaluation
To evaluate the MFSP predictor, we adopt the proposed method with EU as the final model for the MIDR predictor and utilize it to train the MFSP predictor. As described in Section V-B, we consider two training strategies:
  • De novo training of a new network from scratch.
  • TL with GNN by reusing the GNN module, including the initial encoders, from the MIDR predictor.
@@
  • MSE loss focuses on the accuracy of predicted coordinates.
  • Hybrid loss combines the MSE loss with the MIDR-based loss for better ranking consistency.
The network architecture of the MFSP predictor is almost identical to that of the MIDR predictor, consisting of an initial encoder, a GNN, pooling layers, and an MLP task head. There are only a few differences as follows:
  • Final MLP head: The task head in the MFSP predictor replaces the linear layer in the MIDR predictor with a single-hidden-layer MLP. The hidden layer’s dimension matches the graph embedding (twice the node embedding dimension), and the output layer has a dimension of 3, representing the predicted coordinates $(x,y,z)$. A “sigmoid” activation function is applied to normalize the predicted coordinates to the range $[0,1]$, aligned with the building normalized spatial dimensions.
  • De Novo training: The initial node input features include only the spatial coordinates of the nodes. Fire point information is excluded, as no prior knowledge from the MIDR predictor is utilized in this case.
  • TL with GNN training: The GNN module and initial encoders from the MIDR predictor are directly reused. Moreover, the fire point information is replaced with a VFP to ensure consistency in the input format. This enables the network to generalize its understanding of the structural information without relying on specific fire locations.
By evaluating these configurations and loss functions, the MFSP predictor aims to efficiently and accurately predict the MFSP within the building at hand.
@@ VII-B1 Evaluation metrics
In Section V-A, we introduce the “pseudo-labeling” method, which generates pseudo ground truth coordinates for the MFSP using the well-trained MIDR predictor. These pseudo-labels are used to evaluate the MFSP predictor through various metrics, capturing both point-level and room-level performances. Specifically, they are generated by selecting the fire point with the highest predicted MIDR for each structure, and subsequently used to train and evaluate the MFSP predictor. This approach allows leveraging the unlabeled data effectively, as the MIDR predictor provides a computationally efficient and accurate surrogate for the FEA simulations.
For a given structure $m$, the distance error between predicted MFSP and the pseudo ground truth MFSP is defined as:
@@
where $\widehat{x}_{m}$, $\widehat{y}_{m}$, $\widehat{z}_{m}$ represent the $x$, $y$, $z$ coordinates predicted by the MFSP predictor, while $x_{m}^{\,\text{pgt}}$, $y_{m}^{\,\text{pgt}}$, $z_{m}^{\,\text{pgt}}$ denote the pseudo ground truth MFSP coordinates obtained using the MIDR predictor. This metric provides a direct measure of the prediction accuracy at the point level. However, as the pseudo ground truth is defined at a room-level granularity, its reliability may vary depending on the building’s size and complexity. For practical evaluations, two room-level metrics are introduced, namely, room distance error and room rank. These metrics treat rooms as the basic unit of measurements, reflecting how well the MFSP predictor identifies the Most Fire-Sensitive Room (MFSR), i.e., the room where the MFSP is located.
Room distance error measures the distance between the predicted MFSR and the pseudo ground truth MFSR, assuming all edges (beams and columns) of the building structure have unit length. For structure $m$, let the room dimensions be $\tilde{x}_{m}$, $\tilde{y}_{m}$, $\tilde{z}_{m}$ for the respective length, width, and height. The coordinates of the center point of the predicted MFSR are denoted as $\left(\widehat{x}_{m}^{\prime},\widehat{y}_{m}^{\prime},\widehat{z}_{m}^{\prime}\right)$. The room distance error is then computed as:
@@
This metric provides a scaled distance error measure, accounting for the room size and the structural dimensions.
Room rank evaluates the ranking of the room containing the predicted MFSP based on the pseudo-labeled MIDR values. For structure $m$ with $N_{m}$ pseudo fire cases, we sort the pseudo-labeled room-center MIDRs in a descending order, and index the rooms by the fire case occurring at each room’s center, i.e., re-number the rooms to satisfy $d_{m,1,\max}^{\,\text{pgt}}>d_{m,2,\max}^{\,\text{pgt}}>\cdots>d_{m,N_{m},\max}^{\,\text{pgt}}$. If the predicted MFSP falls into the $l$-th ranked room, the room rank is accordingly defined as $r_{m}=l$. Lower ranks indicate better prediction performance, as the predicted MFSR with such low rank is closer to the MFSP.
In addition to the distance error, room distance error, and room rank, the predicted MIDR at the identified MFSP can also serve as an evaluation metric. Higher MIDR values reflect the predictor’s ability to identify the MFSPs, directly correlating with the severity of the structural response. These evaluation metrics collectively assess the MFSP predictor’s performance, balancing point-level precision, room-level granularity, and the functional implications of the predictions. The inclusion of room distance error and room rank ensures that the evaluation aligns with practical building-level assessments, where rooms are typically used as reference units for structural analysis and fire safety planning.
TABLE VII: Evaluation results of the MFSP predictor for different training methods and loss functions. Numbers in boldface indicate best results; values in parentheses show performance multiples (indicated with “$\times$”) relative to the reference configuration, i.e., TL with GNN, Hybrid ($w_{1}=50$).
@@
VII-B2 Evaluation results
Using the metrics defined in the previous section, we evaluate the performance of the MFSP predictor by calculating average quantities across all building structures. Since the MIDR predictor provides a computationally efficient agent model to rapidly generate pseudo labels for unlabeled data, the large unlabeled dataset can be effectively utilized for training the MFSP predictor. For this evaluation, the unlabeled dataset is divided into $80\%$ for training and $20\%$ for testing. The evaluation results for various training methods and loss functions are summarized in Table VII. For the “Hybrid loss”, Equation (12), the weight for the MSE term is fixed at $w_{2}=1\,\mathrm{m}^{-2}$, while the weight for the MIDR term is varied as $w_{1}=10$, $50$, or $100$. In Table VII, the TL with GNN, Hybrid results for $w_{1}=50$, using both small and large networks, are taken as the reference cases for the listed obtained multipliers $\mu$ in parentheses, i.e., $(\mu\times)$.